# Asymptotically conformal CFL quark matter within a nonlocal chiral quark model

### v1

@@ Abstract
We propose a three-flavor nonlocal NJL model of quark matter with the scalar attractive, vector repulsive and diquark pairing interaction channels. The model is treated within the separable approximation to obtain the EoS of cold quark matter. The analysis of the high density asymptotics of the model allows us to conclude about its qualitative agreement with the perturbative QCD. Particularly, a color superconducting CFL state is found to be the ground one at high densities. The conformal limit of speed of sound and dimensionless interaction measure are also shown to be reached from below and above, respectively. The model is applied to modelling NSs within the scenario of early quark deconfinement triggered by the gravitational instability of the NS matter due to the BEC of a spin-color-flavor singlet three-diquark bound state, the light sexaquark, stable against the weak and strong decays. A conservative estimate of the sexaquark mass based on the most up-to-date results from QCD sum rules and a constituent quark model assumes an early quark deconfinement with the onset mass below $1~{}\rm M_{\odot}$. The scenario is shown to be consistent with the present observational constraints on the mass-radius relation and tidal deformability of NSs. Given the fact that the proposed model by construction exhibits an asymptotically conformal behavior, we consider the question about approximately conformal quark matter in NSs. We report a non-perturbative energy density range, which is inaccessible in NSs and where the speed of sound and dimensionless interaction measure of the CFL quark matter simultaneously attain the conformal values.
sexaquarks — quark matter — color superconductivity
@@ I Introduction
Gauge invariance of quantum chromodynamics (QCD) implies an order-by-order agreement of the perturbative calculations within the Coulomb and Lorentz-covariant gauges Zwanziger (1998). At the same time the Coulomb gauge has an advantage of dealing exclusively with the physical degrees of freedom. This becomes crucial at the non-perturbative regimes where the mentioned agreement is spoiled by the interactions of the longitudinal gluons. The Coulomb gauge also provides a renormalization-group invariance of the time-time component of the gluon propagator, thus, making it a physically relevant object. The time-time gluon propagator is also responsible for generating a long-range confining interaction between quarks Gribov (1978); Zwanziger (1997) and simultaneously gives the access to the quasiparticle representation of QCD Szczepaniak and Swanson (2001). This inspired various non-perturbative methods for the low-energy dynamics of quarks including Schwinger-Dyson equations Kocić (1986); Roberts and Williams (1994); Alkofer and von Smekal (2001); Watson and Reinhardt (2007); Binosi and Papavassiliou (2009) and effective Lagrangian approahes rooted in the Nambu–Jona-Lasinio (NJL) model for quark matter Vogl and Weise (1991); Klevansky (1992); Hatsuda and Kunihiro (1994); Buballa (2005).
@@
Unlike the effective quark-quark interactions truncated to the processes mediated by the time-time gluon propagator (see e.g. Ref. Szymański et al. (2023)), the NJL-type approaches extend beyond the vector interaction kernel in agreement with the analysis of spin splittings in heavy quarkonia Schnitzer (1975). Fierz transformation of the (massive) vector boson exchange mode also assumes existence of various channels of the effective quark-quark interaction Buballa (2005). This increases the phenomenological attractivity of the NJL model, which has been serving as a benchmark for modelling the dynamics of quark matter at the regimes typical for heavy ion collisions Vogl and Weise (1991); Klevansky (1992); Hatsuda and Kunihiro (1994); General et al. (2001); Buballa (2005); Gomez Dumm and Scoccola (2002, 2005); Gomez Dumm et al. (2006); Ratti et al. (2007); Câmara Pereira et al. (2020); Maslov and Blaschke (2023) and interiors of neutron stars (NSs) Buballa (2005); Duhau et al. (2004); de Carvalho et al. (2015); Câmara Pereira et al. (2016); Baym et al. (2018); Alvarez-Castillo et al. (2019); Song et al. (2019); Ivanytskyi et al. (2019); Contrera et al. (2022); Blaschke et al. (2023a); Carlomagno et al. (2024).
@@
However, the NJL model suffers a conceptual problem related to its non-renormalizability manifested through a momentum cut off defining its applicability range. This drawback is connected to the contact character of the quark-quark interactions. At the same time, the lattice formulation of QCD (lQCD) indicates that quark interactions act over a certain range in the momentum space Becirevic et al. (2000); Skullerud and Williams (2001); Skullerud et al. (2001); Parappilly et al. (2006); Blossier et al. (2011); Burgio et al. (2012); Oliveira et al. (2019). This implies an additive momentum-dependent mass renormalization in the quark propagator Skullerud et al. (2001), which is also reported within the Hamiltonian formalism of QCD in Coloumb gauge Zwanziger (1997, 1998); Szczepaniak and Swanson (2001); Reinhardt et al. (2018); Quandt et al. (2018). This, in turn, assumes a space-time nonlocality of the effective quark interactions Blaschke et al. (1995); General et al. (2001); Gomez Dumm and Scoccola (2002, 2005); Duhau et al. (2004); Gomez Dumm et al. (2006); Hell et al. (2010). Such a nonlocality also arises within the instanton liquid model Schäfer and Shuryak (1998) and the Schwinger-Dyson approach Kocić (1986); Roberts and Williams (1994); Alkofer and von Smekal (2001); Watson and Reinhardt (2007); Binosi and Papavassiliou (2009). Thus, nonlocality, on one hand, is an important feature of the quark-quark interactions and, on the other one, naturally solves the problem of non-renormalizability of the NJL model Langfeld et al. (1996).
@@
In this work we suggest a version of the three-flavor nonlocal NJL model, which includes the interaction channels omitted before. While spontaneous breaking of chiral symmetry of low-energy QCD is incorporated to the nonlocal NJL model by connecting the dynamically generated mass gap to the chiral condensate Scarpettini et al. (2004); Blaschke et al. (2008); Hell et al. (2009, 2010), application of the model to the regime of high baryon densities requires accounting for the vector repulsive and diquark pairing interactions Buballa (2005); Baym et al. (2018); Song et al. (2019). This question has been partially addressed within the two-flavor nonlocal NJL model, however, either ignoring the diquark pairing channel Plant and Birse (1998) or treating the vector repulsion locally Contrera et al. (2022); Carlomagno et al. (2024, 2024). The previous versions the three-flavor nonlocal NJL model de Carvalho et al. (2015); Blaschke et al. (2023a) have the same drawbacks cured in this work. Thus, we present a nonlocal three-flavor NJL model with vector repulsion and diquark pairing. We also apply the model to modelling the color-superconducting color-flavor-locked (CFL) phase of quark matter Alford et al. (1998, 1999); Schäfer and Wilczek (1999); Alford et al. (2008), which is expected to be the ground state of dense QCD Schäfer and Wilczek (1999). For the sake of simplicity we consider the case when light and strange quarks are degenerated in mass. This corresponds to the so-called CFLL quark matter Blaschke et al. (2023b).
@@
Quark deconfinement at high densities is inevitable due to the asymptotic freedom of QCD, which unbinds hadrons Gross and Wilczek (1973); Politzer (1973). But the open question is whether the deconfinement onsets at the densities reached in the interiors of NSs Baym et al. (2018). Recent constraints on their maximum mass well above the two solar ones along with the corresponding radius Miller et al. (2021); Riley et al. (2021) exceeding the one of a canonical NS Dietrich et al. (2020) are challenging for description within purely hadronic scenarios. In the presence of hyperons these scenarios also come into a tension with the reported high mass of the black widow pulsar PSR J0952-0607 Romani et al. (2022), which can be naturally explained as a NS with a color-superconducting quark matter core Gärtlein et al. (2023). Formation of such cores in the NS interiors generates a characteristic back-bending of the mass-radius relation Alvarez-Castillo et al. (2019); Contrera et al. (2022); Ivanytskyi and Blaschke (2022a, b); Carlomagno et al. (2024); Blaschke et al. (2023b); Gärtlein et al. (2023); Li et al. (2024) providing a good agreement with the mentioned observational constraints. Such back-bending also makes the scenario of hybrid quark-hadron NSs consistent with the enigmatic HESS J1731-347 compact object Sagun et al. (2023). The model agnostic statistical analysis of the observational constraints on the NS mass-radius diagram also assumes existence of quark matter in their interiors Annala et al. (2020, 2023). Finally, quark deconfinement during the supernova explosion suggests a unique mechanism of producing NSs with masses above the two solar ones Fischer et al. (2018).
@@
Weakening the quark-quark interactions in the perturbative domain drives the QCD matter toward the conformal limit Kurkela et al. (2010); Fraga et al. (2014); Gorda et al. (2018). This stimulated intense discussions of the possible nearly conformal behavior of the NS matter as a signature of quark deconfinement Fujimoto et al. (2022); Altiparmak et al. ; Marczenko et al. ; Brandes et al. (2023, 2023); Takatsy et al. (2023); Annala et al. (2023). The nonlocal NJL model presented below is asymptotically conformal by construction, which makes it a great tool for studying this possibility. Below we address this question and demonstrate the existence of a non-perturbative range of densities inaccessible in NSs, where the speed of sound and dimensionless interaction measure of cold quark matter simultaneously attain nearly conformal values.
This paper considers the picture of an early quark deconfinement, which has been argued to be preferred by the observational data Alvarez-Castillo et al. (2019); Alvarez-Castillo (2021); Ivanytskyi and Blaschke (2022a, b, c); Blaschke et al. (2023a, 2022); Contrera et al. (2022); Sagun et al. (2023); Gärtlein et al. (2023). One of its possible reasons corresponds to the recently considered Bose-Einstein condensation (BEC) of a color-spin-flavor singlet sexaqauark state triggering a mechanical instability of hadron matter against the gravitational compression in NSs Blaschke et al. (2022); Shahrbaf et al. (2022). While its spin-3 electrically charged non-strange analogue $d^{*}(2380)$ has been observed experimentally Adlarson et al. (2011), detection of an electrically neutral sexaquark remains a challenge even despite a smaller mass. QCD sum rules Azizi et al. (2020) and a constituent quark model Buccella (2020) predict this mass to be below the strong and weak decay thresholds. This supports the possibility of existence of the BEC of sexaquarks in the baryon rich NS matter. Being unable to resist the gravitational compression in NSs, such a BEC state increases its density until the sexaquarks and other hadrons get dissociated. This causes quark deconfinement Since the flavor content of sexaquarks ($uuddss$) is the same as the one of the diquark triplet ($ud$, $ds$, $su$), the attained phase is the CFL quark matter constituted by the mentioned diquark correlations. As shown below, sexaquark masses below the weak decay threshold assume an early quark deconfinement in NSs lighter than one solar mass. A phenomenological conclusion of this is that all the observed NSs have a quark core. It is worth mentioning, this conclusion is also supported by the recent study of the HESS J1731-347 nature Sagun et al. (2023).
@@ II Nonlocal NJL model for CFLL quark matter
The present consideration is limited to the interaction channels, which acount for the most important aspects of the phenomenology of cold and dense quark matter. Dynamical restoration and breaking of chiral symmetry and generation of the quark mass gap is among them. The minimal scheme for accounting for this aspect requires the scalar interaction channel, which includes eight scalar current densities labelled by the index $a=0,1$
@@
The repulsive interaction in dense quark matter motivated by the non-perturbative gluon exchange Song et al. (2019) is needed for stiffness of quark equation of state (EoS). This stiffness provides a positive feedback to the two solar mass limit problem for NSs Antoniadis et al. (2013); Cromartie et al. (2019); Miller et al. (2021); Riley et al. (2021). Within the nonlocal NJL model such a repulsion is generated by the vector interaction channel given in terms of the vector current density
@@
Any Fermi system with an arbitrary weak attractive interaction at sufficiently high densities is a subject to a Cooper instability Cooper (1956). In dense quark matter this leads to formation of a color-superconducting state consituted by the BEC of diquarks Alford et al. (1998, 1999); Schäfer and Wilczek (1999); Alford et al. (2008). In order to account for this we introduce to the consideration the diquark current density
@@
Here $\lambda_{b}$ are Gell-Mann matrices related to the generators of the SU(3) color group, the indexes $a,b=2,5,7$ correspond to the antisymmetric color-flavor operators, the charge conjugate quark field is $q^{c}=i\gamma_{2}\gamma_{0}\overline{q}^{T}$  [^1] An extended discussion of the diquark operators can be found in the review Buballa (2005).. It is worth noting, that the current (3) corresponds to condensation of scalar diquarks, which are favored by QCD Hess et al. (1998); Alexandrou et al. (2006); Babich et al. (2007); DeGrand et al. (2008), while the axial-vector realization of the diquark ground state (see e.g. Ref. Anselmino et al. (1993)) is not considered in this paper.
@@
Here $m$ is the current quark mass. For the CFLL quark matter considered in this work, it coincides for three quark flavors. In the case of such a mass degeneration densities of all the quark flavors coincide, while color and electric charge neutrality is provided by a common quark chemical potential $\mu$ Blaschke et al. (2023a). The $\beta$-equilibrium is also authomatically provided in this case. The quark chemical potential enters the Lagrangian (4) via the term including the operator $\overline{q}\gamma_{0}q$ of conserved quark number density. The scalar, vector and diquark couplings $G_{S}$, $G_{V}$ and $G_{D}$ control the strength of interaction in the scalar, vector and diquark channels discussed above. The conventional factor $3$ in the last term in Eq. (4) is introduced for the sake of further convenience. Following the commonly accepted practice, in this work we parameterize the vector and diquark couplings via the dimensionless parameters $\eta_{V}=G_{V}/G_{S}$ and $\eta_{D}=G_{D}/G_{S}$.
@@
From now on we want to sharpen the consideration on the MFA. In this case functional integrations over the bosonic fields in Eq. (5) are suppressed, while the fields themselves are replaced by their expectation values. The later can be found by averaging Eqs. (7)-(9). The scalar currents generated by the non-diagonal flavor Gell-Mann matrices do not survive the averaging procedure due to the vanishing of the corresponding functional integrals over the quark fields. This conclusion can be independently drawn from the condition of charge conservation provided by the commutation of the mean-field Lagrangian with the charge matrix $\hat{Q}={\rm diag}(2/3,-1/3,-1/3)$ Hell et al. (2010). This requires the absence of the non-diagonal flavor group generators in the mean-field Lagrangian, i.e. vanishing of the corresponding scalar currents. Furthermore, flavor degeneration in the CFLL quark matter along with the traceless character of $\tau_{3}$ and $\tau_{8}$ leads to vanishing of the expectation values of the scalar current densities $s_{3}$ and $s_{8}$. Consequently, only the scalar field $\sigma_{0}$ attains a non-vanishing value within the MFA to the present model, i.e. $\langle\sigma_{a}\rangle=\sigma\delta_{a0}$ with
@@
Locking of the color and flavor indexes of quarks in the CFL matter means that only three diquark condensates survive the averaging in the ground state Pisarski and Rischke (1999). It is possible to define a color-flavor rotation under which they are the condensates generated by the antisymmetric SU(3) flavor and color generators with $a=b$ Buballa (2005). The same holds for the diquark current densities since they have the same color-flavor structure as the diquark condensates, i.e. $\langle d_{ab}\rangle=\langle d_{aa}\rangle\delta_{ab}$. In the CFLL phase all the quark color-flavor states are equivalent making three non-vanishing diquark condensates equal. This allows us to introduce
@@
Evaluating the quark term in Eq. (13) involves solving the single particle energies of quarks shifted by the quark chemical potential. They are defined as a solution of the equation $\det\mathcal{S}^{-1}=0$ with respect to the temporal component of the quark four-momentum. In the general case of different current quark masses finding this solution is only possible numerically, while flavor degeneration in the CFLL quark matter allows to do it analytically. This is equivalent to diagonalizing the Nambu-Gorkov propagator. The corresponding basis includes singlet and octet states of paired quarks Shovkovy and Wijewardhana (1999). In what follows these states are labelled by the subscript index $j=\rm sing$ and $j=\rm oct$, while their degeneracy factors are $d_{\rm sing}=2\times 1$ and $d_{\rm oct}=2\times 8$, respectively, where the factors $2$ account for the the spin degrees of freedom. Thus, the single particle energies of quarks shifted by the chemical potential are

### v2

@@ Abstract
We propose a three-flavor nonlocal NJL model of quark matter with the scalar attractive, vector repulsive and diquark pairing interaction channels. The model is treated within the separable approximation to obtain the EoS of cold quark matter. The analysis of the high density asymptotics of the model allows us to conclude about its qualitative agreement with the perturbative QCD. Particularly, a color superconducting CFL state is found to be the ground one at asymptotically high densities. The conformal limit of speed of sound and dimensionless interaction measure are also shown to be reached from below and above, respectively. The model is applied to modelling NSs within the scenario of early quark deconfinement triggered by the gravitational instability of the NS matter due to the BEC of a spin-color-flavor singlet three-diquark bound state, the light sexaquark, stable against the weak and strong decays. A conservative estimate of the sexaquark mass based on the most up-to-date results from QCD sum rules and a constituent quark model implies an early quark deconfinement with the onset mass below $1~{}\rm M_{\odot}$. The scenario is shown to be consistent with the present observational constraints on the mass-radius relation and tidal deformability of NSs. Given the fact that the proposed model by construction exhibits an asymptotically conformal behavior, we consider the question about approximately conformal quark matter in NSs. We report a non-perturbative energy density range, which is inaccessible in NSs and where the speed of sound and dimensionless interaction measure of the CFL quark matter simultaneously attain the conformal values.
quark matter — color superconductivity — conformal limit
@@ I Introduction
Gauge invariance of quantum chromodynamics (QCD) implies an order-by-order agreement of the perturbative calculations within the Coulomb and Lorentz-covariant gauges Zwanziger (1998). At the same time the Coulomb gauge has an advantage of dealing exclusively with the physical degrees of freedom. This becomes crucial at the non-perturbative regimes where the mentioned agreement is spoiled by the interactions of the longitudinal gluons. The Coulomb gauge also provides a renormalization-group invariance of the time-time component of the gluon propagator, thus, making it a physically relevant object. The time-time gluon propagator is also responsible for generating a long-range confining interactions between quarks Gribov (1978); Zwanziger (1997). Within the Coulomb gauge it simultaneously allows constructing a non-perturbative Hamiltonian, which can be expanded in a quasiparticle basis spaning the Fock space Szczepaniak and Swanson (2001). This gives the access to the quasiparticle representation of QCD and justifies various non-perturbative methods for the low-energy dynamics of quarks including Schwinger-Dyson equations Kocić (1986); Roberts and Williams (1994); Alkofer and von Smekal (2001); Watson and Reinhardt (2007); Binosi and Papavassiliou (2009) and effective Lagrangian approahes rooted in the Nambu–Jona-Lasinio (NJL) model for quark matter Vogl and Weise (1991); Klevansky (1992); Hatsuda and Kunihiro (1994); Buballa (2005).
@@
Unlike the effective quark-quark interactions truncated to the processes mediated by the time-time gluon propagator (see e.g. Ref. Szymański et al. (2023)), the NJL-type approaches extend beyond the vector interaction kernel in agreement with the analysis of spin splittings in heavy quarkonia Schnitzer (1975). Fierz transformation of the (massive) vector boson exchange mode also suggests an argument in favor of existence of various channels of the effective quark-quark interaction Buballa (2005). This makes the NJL model phenomenologically attractive for modelling the dynamics of quark matter at the regimes typical for heavy ion collisions Vogl and Weise (1991); Klevansky (1992); Hatsuda and Kunihiro (1994); General et al. (2001); Buballa (2005); Gomez Dumm and Scoccola (2002, 2005); Gomez Dumm et al. (2006); Ratti et al. (2007); Câmara Pereira et al. (2020); Maslov and Blaschke (2023) and interiors of neutron stars (NSs) Buballa (2005); Duhau et al. (2004); de Carvalho et al. (2015); Câmara Pereira et al. (2016); Baym et al. (2018); Alvarez-Castillo et al. (2019); Song et al. (2019); Ivanytskyi et al. (2019); Contrera et al. (2022); Blaschke et al. (2023a); Carlomagno et al. (2024).
@@
However, the NJL model suffers a conceptual problem related to its non-renormalizability manifested through a momentum cut off defining its applicability range. This drawback is connected to the contact character of the quark-quark interactions. At the same time, the lattice formulation of QCD (lQCD) indicates that quark interactions act over a certain range in the momentum space Becirevic et al. (2000); Skullerud and Williams (2001); Skullerud et al. (2001); Parappilly et al. (2006); Blossier et al. (2011); Burgio et al. (2012); Oliveira et al. (2019). This implies an additive momentum-dependent mass renormalization in the quark propagator Skullerud et al. (2001), which is also reported within the Hamiltonian formalism of QCD in Coloumb gauge Zwanziger (1997, 1998); Szczepaniak and Swanson (2001); Reinhardt et al. (2018); Quandt et al. (2018). This, in turn, means a space-time nonlocality of the effective quark interactions Blaschke et al. (1995); General et al. (2001); Gomez Dumm and Scoccola (2002, 2005); Duhau et al. (2004); Gomez Dumm et al. (2006); Hell et al. (2010). Such a nonlocality also arises within the instanton liquid model Schäfer and Shuryak (1998) and the Schwinger-Dyson approach Kocić (1986); Roberts and Williams (1994); Alkofer and von Smekal (2001); Watson and Reinhardt (2007); Binosi and Papavassiliou (2009). Thus, nonlocality, is an important feature of the quark-quark interactions that should be accounted for within the effective approaches of the NJL-type by introducing non-contact correlations between quark currents. Remarkably, this naturally solves the problem of non-renormalizability of the NJL model, since the momentum cutoff mentioned above is not required in this case Langfeld et al. (1996).
@@
While spontaneous breaking of chiral symmetry of low-energy QCD is incorporated to the nonlocal NJL model by connecting the dynamically generated mass gap to the chiral condensate Scarpettini et al. (2004); Blaschke et al. (2008); Hell et al. (2009, 2010), application of the model to the regime of high baryon densities requires accounting for the vector repulsive and diquark pairing interactions Buballa (2005); Baym et al. (2018); Song et al. (2019). The last two interaction channels were considered within the two-flavor nonlocal NJL model, however, with a drawback of either ignoring the diquark pairing Plant and Birse (1998) or treating the vector repulsion locally Contrera et al. (2022); Carlomagno et al. (2024, 2024). The previous versions the three-flavor nonlocal NJL model de Carvalho et al. (2015); Blaschke et al. (2023a) have the same drawbacks cured in this work. We present a nonlocal three-flavor NJL model with vector repulsion and diquark pairing. We also apply it to modelling the color-superconducting color-flavor-locked (CFL) phase of quark matter Alford et al. (1998, 1999a); Schäfer and Wilczek (1999a); Alford et al. (2008), which is expected to be the ground state of dense QCD Schäfer and Wilczek (1999a); Son (1999). For the sake of simplicity we consider the case when light and strange quarks are degenerated in mass. This corresponds to the so-called CFLL quark matter Blaschke et al. (2023b). Based on a simple kinematic criterion of ”unlocking” strange quarks from the CFL phase if their mass is high enough Schäfer and Wilczek (1999b); Alford et al. (1999b); Abuki (2003), which corresponds to the two-flavor color-superconductivity in quark matter (2SC) Buballa (2005), we argue that such unlocking does not happen for the considered strengths of the diquark pairing interaction and that the CFLL picture describes dense QCD matter with a reasonable accuracy. Later we use the developed approach to model NSs with color-superconducting quark cores. It is worth mentioning that quark pairing beyond the 2SC and CFL patterns has been shown to be important for the phenomenology of these astrophysical objects Grigorian et al. (2005); Popov et al. (2006); Kojo et al. (2021). The color-spin locking (CSL) suggests a mechanism of a single-flavor color-superconductivity, which is able to explain such pairing. The CSL mechanism is applicable even in the case of a substantial Fermi momentum mismatch among light and strange quark flavors caused by the large mass of the latter Schäfer (2000); Schmitt et al. (2003); Marhauser et al. (2007); Aguilera et al. (2006); Feng et al. (2010).
@@
Quark deconfinement at high densities is inevitable due to the asymptotic freedom of QCD, which unbinds hadrons Gross and Wilczek (1973); Politzer (1973). But the open questions are whether the deconfinement onsets at the densities reached in the interiors of NSs and what is the corresponding onset mass Baym et al. (2018)? In the case when the latter is below one solar mass we refer to an early quark deconfinement in NSs. Recent constraints on their maximum mass well above $2{\rm M}_{\odot}$ along with the corresponding radius Miller et al. (2021); Riley et al. (2021) exceeding the one of a canonical NS with the mass $1.4{\rm M}_{\odot}$ Dietrich et al. (2020) are challenging for description within purely hadronic scenarios. In the presence of hyperons these scenarios also come into a tension with the reported high mass of the black widow pulsar PSR J0952-0607 Romani et al. (2022), which potentially provide a more stringent constraint on the NS maximum mass that, nevertheless, should be taken cautiously due to the unknown angular momentum of this accreting system. On the other hand, this pulsar can be naturally explained as a NS with a color-superconducting quark matter core Gärtlein et al. (2023). Formation of such cores in the NS interiors generates a characteristic back-bending of the mass-radius relation Alvarez-Castillo et al. (2019); Contrera et al. (2022); Ivanytskyi and Blaschke (2022a, b); Carlomagno et al. (2024); Blaschke et al. (2023b); Gärtlein et al. (2023); Li et al. (2024) providing a good agreement with the mentioned observational constraints. Such back-bending also leads to relatively small radii of NSs with mass about $1\rm M_{\odot}$. This makes the scenario of hybrid quark-hadron NSs consistent with the enigmatic HESS J1731-347 compact object Sagun et al. (2023). The model agnostic statistical analysis of the observational constraints on the NS mass-radius diagram also reports existence of quark matter in their interiors Annala et al. (2020, 2023). Finally, quark deconfinement during the supernova explosion suggests a unique mechanism of producing NSs with masses above the two solar ones Fischer et al. (2018). Based on this, the paper considers the scenario of an early quark deconfinement, which has been argued to be prefered by the observational data Alvarez-Castillo et al. (2019); Alvarez-Castillo (2021); Ivanytskyi and Blaschke (2022a, b, c); Blaschke et al. (2023a, 2022); Contrera et al. (2022); Sagun et al. (2023); Gärtlein et al. (2023). A phenomenological consequence of this is that all the observed NSs have a quark core. It is worth mentioning, this possibility is also supported by the recent study of the HESS J1731-347 nature Sagun et al. (2023). As a possible trigger of an early deconfinement in NSs, we discuss the recently proposed Bose-Einstein condensation (BEC) of a color-spin-flavor singlet sexaqauark state causing a mechanical instability of hadron matter against the gravitational compression in NSs and onset of the CFL quark matter Blaschke et al. (2022); Shahrbaf et al. (2022), which. Nevertheless, thim mechsnism is not build into the model.
@@
Weakening the quark-quark interactions in the perturbative domain drives the QCD matter toward the conformal limit Kurkela et al. (2010); Fraga et al. (2014); Gorda et al. (2018); Fernandez and Kneur (2022). This stimulated intense discussions of the possible nearly conformal behavior of the NS matter as a signature of quark deconfinement Fujimoto et al. (2022); Altiparmak et al. ; Marczenko et al. ; Brandes et al. (2023, 2023); Takatsy et al. (2023); Annala et al. (2023). The nonlocal NJL model presented below is asymptotically conformal by construction. This enables using it to test how close to the conformal behavior is the QCD matter at the densities reached in the centers of the heaviest NSs. Below we address this question and demonstrate the existence of a non-perturbative range of densities inaccessible in NSs, where the speed of sound and dimensionless interaction measure of cold quark matter simultaneously attain nearly conformal values.
@@ II Nonlocal NJL model for CFLL quark matter
The present consideration is limited to the interaction channels, which acount for the most important aspects of the phenomenology of cold and dense quark matter. Dynamical restoration and breaking of chiral symmetry and generation of the quark mass gap is among them. The minimal scheme for accounting for this aspect requires the scalar interaction channel, which includes eight scalar current densities labeled by the index $a=0,8$
@@
The repulsive interaction in dense quark matter motivated by the non-perturbative gluon exchange Song et al. (2019) is needed for stiffness of quark equation of state (EoS). This enables NSs with quark cores to reach two solar masses Antoniadis et al. (2013); Cromartie et al. (2019); Miller et al. (2021); Riley et al. (2021). Within the nonlocal NJL model such a repulsion is generated by the vector interaction channel given in terms of the vector current density
@@
Any Fermi system with an arbitrary weak attractive interaction at sufficiently high densities is a subject to a Cooper instability Cooper (1956). In dense quark matter this leads to formation of a color-superconducting state consituted by the BEC of diquarks Alford et al. (1998, 1999a); Schäfer and Wilczek (1999a); Alford et al. (2008). In order to account for this we introduce to the consideration the diquark current density
@@
Here $\lambda_{b}$ are Gell-Mann matrices related to the generators of the SU(3) color group, the indexes $a,b=2,5,7$ correspond to the antisymmetric color-flavor operators, the charge conjugate quark field is $q^{c}=i\gamma_{2}\gamma_{0}\overline{q}^{T}$  [^1] An extended discussion of the diquark operators can be found in the review Buballa (2005).. It is worth noting that the current (3) corresponds to condensation of scalar diquarks, which are favored by QCD Hess et al. (1998); Alexandrou et al. (2006); Babich et al. (2007); DeGrand et al. (2008), while the axial-vector realization of the diquark ground state (see e.g. Ref. Anselmino et al. (1993)) is not considered in this paper.
@@
Here $m$ is the current quark mass. For the CFLL quark matter considered in this work this mass is the same for three quark flavors. In the case of such a mass degeneration densities of all the quark flavors coincide, while color and electric charge neutrality is provided by a common quark chemical potential $\mu$ Blaschke et al. (2023a). The $\beta$-equilibrium is also authomatically provided in this case. The quark chemical potential enters the Lagrangian (4) via the term including the operator $\overline{q}\gamma_{0}q$ of conserved quark number density. The scalar, vector and diquark couplings $G_{S}$, $G_{V}$ and $G_{D}$ control the strength of interaction in the scalar, vector and diquark channels discussed above. The conventional factor $3$ in the last term in Eq. (4) is introduced for the sake of further convenience. Following the commonly accepted practice, in this work we parameterize the vector and diquark couplings via the dimensionless parameters $\eta_{V}=G_{V}/G_{S}$ and $\eta_{D}=G_{D}/G_{S}$.
@@
From now on we want to sharpen the consideration on the MFA. In this case functional integrations over the bosonic fields in Eq. (5) are suppressed, while the fields themselves are replaced by their expectation values. The latter can be found by averaging Eqs. (7)-(9). The scalar currents generated by the non-diagonal flavor Gell-Mann matrices do not survive the averaging procedure due to the vanishing of the corresponding functional integrals over the quark fields. This conclusion can be independently drawn from the condition of charge conservation provided by the commutation of the mean-field Lagrangian with the charge matrix $\hat{Q}={\rm diag}(2/3,-1/3,-1/3)$ Hell et al. (2010). This requires the absence of the non-diagonal flavor group generators in the mean-field Lagrangian, i.e. vanishing of the corresponding scalar currents. Furthermore, flavor degeneration in the CFLL quark matter along with the traceless character of $\tau_{3}$ and $\tau_{8}$ leads to vanishing of the expectation values of the scalar current densities $s_{3}$ and $s_{8}$. Consequently, only the scalar field $\sigma_{0}$ attains a non-vanishing value within the MFA to the present model, i.e. $\langle\sigma_{a}\rangle=\sigma\delta_{a0}$ with
@@
Locking of the color and flavor indexes of quarks in the CFL matter means that only three diquark condensates survive the averaging in the ground state Pisarski and Rischke (1999). It is possible to define a color-flavor rotation under which they are the condensates generated by the antisymmetric SU(3) flavor and color generators with $a=b$ Buballa (2005). The same holds for the diquark current densities since they have identical color-flavor structure as the diquark condensates, i.e. $\langle d_{ab}\rangle=\langle d_{aa}\rangle\delta_{ab}$. In the CFLL phase all the quark color-flavor states are equivalent making three non-vanishing diquark condensates equal. This allows us to introduce
@@
Evaluating the quark term in Eq. (13) involves solving the single particle energies of quarks shifted by the quark chemical potential. They are defined as a solution of the equation $\det\mathcal{S}^{-1}=0$ with respect to the temporal component of the quark four-momentum. In the general case of different current quark masses finding this solution is only possible numerically, while flavor degeneration in the CFLL quark matter allows to do it analytically. This is equivalent to diagonalizing the Nambu-Gorkov propagator. The corresponding basis includes singlet and octet states of paired quarks Shovkovy and Wijewardhana (1999). In what follows these states are labeled by the subscript index $j=\rm sing$ and $j=\rm oct$, while their degeneracy factors are $d_{\rm sing}=2\times 1$ and $d_{\rm oct}=2\times 8$, respectively, where the factors $2$ account for the the spin degrees of freedom. Thus, the single particle energies of quarks shifted by the chemical potential are